A 46.0-kg penguin on frictionless ice holds onto the end of a horizontal spring with spring constant 170 N/m and oscillates back and forth. What is its period of oscillation? Answer in seconds.
1
Expert's answer
2011-02-24T05:52:36-0500
According to the mass-period expression:
T = 2π √ (m/k) = 2*3.14 √ (46/170) = 3.26 s